OpenStudy (anonymous):

c^2-2c/6c-2 diveded by 2-c/1-3c?

hero (hero):

step by step?

OpenStudy (anonymous):

yes please

hero (hero):

Well first off the answer is c/2 and I'll show you why

OpenStudy (anonymous):

c^2 - 2c / 6c - 2 x 1 - 3c/2 -c (numerators) [-c(2 - c)] x (1 -3c) (denominators) [-2(1-3c)] x (2 -c) now (2 - c) will be canceled and so will be (1 - 3c) we'll be left with -c in numerator and -2 in denominator making the fraction -c/-2 or c/2

OpenStudy (anonymous):

c^2 - 2c = c(c-2) = -c (-c + 2) = -c(2-c) and 6c -2 = 2 (3c - 1) = -2 (-3c+1) = -2 (1-3c) and then the rest

hero (hero):

\[\frac{c^2-2c}{6c-2} \div \frac{2-c}{1-3c} = \frac{c^2-2c}{6c-2} \times \frac{1-3c}{2-c}=\frac{c(c-2)}{2(3c-1)} \times \frac{3c -1}{c-2}=\frac{c}{2}\]